OpenStudy (foreverandalways13):

USE SUBSTITUTION... SOLVE FOR X y=x-2 y=-2x+7

OpenStudy (foreverandalways13):

HELP!!!!!!!!

OpenStudy (marissalovescats):

Well y=x-2

OpenStudy (marissalovescats):

That means: x-2=-2x+7

OpenStudy (marissalovescats):

See how I did that? @foreverandalways13

OpenStudy (foreverandalways13):

yeah i think so

OpenStudy (marissalovescats):

because it already gave you a y= you can use substitution to plug in x-2 is y because hey. it says y=x-2. Just like if it said y=3 and use substitution, you'd plug 3 in for y. Understand?

OpenStudy (marissalovescats):

So what would x equal? In x-2=-2x+7 @foreverandalways13

OpenStudy (foreverandalways13):

-5

OpenStudy (foreverandalways13):

right?

OpenStudy (foreverandalways13):

@marissalovescats

OpenStudy (anonymous):

So x-2=-2x+7, add 2 to both sides, so x=-2x+9, subtract one x from both sides, so that -3x+9=0, subtract 9 from both sides; -3x = -9. What is x equal to?

OpenStudy (foreverandalways13):

3

OpenStudy (marissalovescats):

3 yup! Great job!! :)

OpenStudy (anonymous):

Well done :)

OpenStudy (foreverandalways13):

thanks guys!!!

OpenStudy (anonymous):

Any time